Weights of hummingbirds are normally distributed with a mean of 48 grams and a standard deviation of 3.5 grams. What is the probability that a randomly selected hummingbird weighs between 44 and 52 grams
Answer by Boreal(15235)

it is a z between x and y
z=(44-48)/3.5=-4/3.5=-1.14
z=(52-48)/3.5+4/3.5=+1.14
probability -1.14< z < 1.14 is 0.7457